OpenStudy (anonymous):

solve for the variable ln(x+2)-ln(x-1)=e

OpenStudy (anonymous):

sorry the question is =1 not=e ln(x+2)-ln(x+1)=1

OpenStudy (anonymous):

I used the a=log(base B)(b^a) to do ln(x+2)/(x-1)=ln(e)

OpenStudy (anonymous):

but I don't know how to do x+2/x-1

OpenStudy (anonymous):

Am I doing the problem correctly as of right now?

OpenStudy (anonymous):

\[\ln(x+2) - \ln(x-1) = 1 \implies \ln(\frac{ x+2 }{ x-1 }) = 1\] From here, I use the property that \(log_{b}m = x \implies\ b^{x} = m\) The base of ln is e, so I can turn this into: \[\ln(\frac{ x+2 }{ x-1 }) = 1 \implies e = \frac{ x+2 }{ x-1 }\]

OpenStudy (anonymous):

Ok... so then do we multiply both sides by x-1? ex-e=x+2

OpenStudy (anonymous):

Right. And then proceed to solve for x.

OpenStudy (anonymous):

... sorry I don't have much experience with e so is it just going to be ex-x=e+2

OpenStudy (anonymous):

e is just a number, it just looks weird at first. But thats right. Now you can just factor x out of the left two terms.

OpenStudy (anonymous):

wait..... won't it work as well if I did ln(x+2/x-1)=ln(e) using the a=log((log base)^a)?

OpenStudy (anonymous):

wait..... nvm sorry I confused myself

OpenStudy (anonymous):

x(e-1)=e+2?

OpenStudy (anonymous):

Yep. And then just divide by e-1 to get \(x = \frac{e+2}{e-1}\)

OpenStudy (anonymous):

Ok I got it now. Thanks!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

but how would it work for ln(x+2)-ln(x-1)=1 x+2-(x-1)=e (e+2/e-1)+2-(e+2/e-1)+1=e (e+2/e-1)-(e+2/e-1)=e-3

OpenStudy (anonymous):

or do we divide the ln() on the left side to get (e+2/e-1)/(e+2/e-1)=e+3

OpenStudy (anonymous):

*e-3 not e+3